OpenStudy (anonymous):

33l^2-l^3-33l Find the zeros of this polynomial function.

OpenStudy (anonymous):

Don't you want to maximize the volume?

OpenStudy (anonymous):

Would I be able to use the quadratic formula?

OpenStudy (anonymous):

No quadratics only work with powers of 2. There are cubic formulas, but they take up alot of room. You could use calculus! lol. The maximum volume is 4.61*10^3 at l=21.5cm. the zeros are: 0 and 32.

OpenStudy (anonymous):

Hm. O.k thank you for the helpful explanation! How about this: Write an inequality that describes the possible values of x for our box problem. Remember that the y value is the volume, and that none of the dimensions can be 0 or negative.

OpenStudy (anonymous):

(This maybe referring back to my other scenerio.)

OpenStudy (anonymous):

3<x<30. h=3-x, so if x is 3 or less, h is zero or negative. w=30-h, so if h is 30 or more, w is zero or negative.

OpenStudy (anonymous):

Volume is positive across all of those values too. You'll have to trust my TI-Nspire CAS for that though.

OpenStudy (anonymous):

Thank you! I don't have the proper calculator handy :/ Do you mind? Use your calculator to find the value of x that gives the maximum volume. Use 2nd TRACE (CALC) and choose the maximum command. Round your answer to the nearest tenth.

OpenStudy (anonymous):

1) TI-Nspires don't work using those exact commands. 2) The maximum is posted in my second post. 3) Because I'm lazy too: V=4.61*10^3 cubic cm at x=21.5cm

OpenStudy (anonymous):

Also, I'm sorry about the too part, it was not meant to insult you in any way.

OpenStudy (anonymous):

Oh lol :) No it's fine! I am! :) Thank you for the help kd8cpk!

OpenStudy (anonymous):

mathsux4real, What is that symbol that looks like a vertical bar? Is it a lower case L?

OpenStudy (anonymous):

Oh yeah... lowercase L=l

OpenStudy (anonymous):

A Mathematica solution with comments is attached.

OpenStudy (anonymous):

Wow THANKS

OpenStudy (anonymous):

The numeric value of the roots are:\[\{\{x\to 0.\},\{x\to 1.03229\},\{x\to 31.9677\}\} \]

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Your welcome.